40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 2 23:11:43 2010, Koji, Update, SUS, ITMX hung 
    Reply  Wed Mar 3 02:57:29 2010, Jenne, Update, SUS, ITMX hung 
    Reply  Fri Mar 5 11:21:18 2010, steve, Update, SUS, used OSEMs are magnetic 
       Reply  Fri Mar 5 18:04:56 2010, rana, Update, SUS, used OSEMs are magnetic 
Message ID: 2651     Entry time: Tue Mar 2 23:11:43 2010     Reply to this: 2652   2658
Author: Koji 
Type: Update 
Category: SUS 
Subject: ITMX hung 

Jenne and Koji

We successfully hung ITMX on the SOS. Side magnet is ~2mm off from the center of the OSEM. ITMX aligned using the QPD. The OSEMs changes the alignment. It looks that something magnetic is inside the OSEM PD or LED.

Reguled ITMY side magnet.

Cleaned up the lab for the safety inspection.

ELOG V3.1.3-